I need help on a few percent problems please.
1) 3 is what percent of 30?
2) 20% of 50 is?
3) 15 is what percent of 75%
4) 30% of 70 is?
Thanks beforehand Joanie

Answer by Mathtut(3670)

1) 3 is what percent of 30?
2) 20% of 50 is?
3) 15 is what percent of 75
4) 30% of 70 is?
:
1)3=x30
:
x=3/30=.1(100)=10%
:
2).2(50)=10
:
3)15=75x
:
x=15/75=.2(100)=20%
:
4).3(70)=21